7 Capstone PrODIEl All Day Couscous Company is a wholesale distributor of candy. The company services grocery, convenience, and drugstores in a large metropolitan area. Small but steady growth in sales has been achieved by the All-Day Couscous Company over the past few years while candy prices have been increasing. The company is formulating its plans for the coming fiscal year July 1 2015 to June 30. 2016. Presented below are the data used to project the current years (July 1, 2015 to June 30 2016) $4.00 per box Average selling price Average variable expenses Direct materials Selling expenses $2.00 per box 0.40 per box $2.40 Total $160,000 280,000 $440,000 Annual fixed expenses Selling Administrative Total $1,560,000 Expected annual sales volume (390,000 boxes) Manufacturers of direct materials have announced that they will increase prices of their products an average of 15 percent in the coming year due to increases in materials and labor expenses. All-Day Couscous Company expects that all other expenses will remain at the same rates or levels as the current year. REQUIRED 1. What is All-Day Couscous Companys breakeven point in boxes of couscous for the current year? What selling price per box must All-Day Couscous Company charge to cover the 15 percent increase in the cost of direct materials and still maintain the current average contribution margin per box? 2. What volume of sales in boxes must the All-Day Couscous Company achieve in the coming year irn order to maintain the same operating income ($184,000) as projected for the current year if the selling price remains at $4.00 per box and the cost of direct materials increases 15 percent? All-Day Couscous is considering adding a new product line, Taco Couscous, to their range of products. They expect that these would sell for $6.00 for a box and would have manufacturing cost of $3.00 have selling costs of $1.00 per box. There would be no additional fixed costs They expect that this new product would be about 1 coming year. What volume of sales in boxes of couscous and Taco must the All-Day Couscous Company achieve in the coming year to maintain the same operating income as projected for the current year if the selling price of couscous remains at $4.00 per box and the cost of direct materials increases 15 percent as in part 3? Briefly explain the reason for the difference in the total volume in parts 3 and 4 4. 0% of their sales volume in boxes in the
Answer #1

Answer 1:

Contribution per box = Average sales price - average variable cost per box = $4.00 - $2.40 = $1.60

Annual fixed expenses = $440,000

Break-even point in boxes = Fixed expenses / contribution per unit = 440000 / 1.60 = 275,000

Break-even point in boxes = 275,000 boxes

Answer 2:

New variable cost per box = $2.00 * (1 + 15%) +$0.40 = $2.70

Selling price required to maintain same average contribution margin per box = Average contribution per box + Increased variable cost per box = $1.60 + $2.70 = $4.30

Selling price required to maintain same average contribution margin per box = $4.30 per box

Answer 3:

Contribution margin per box = $4.00 - $2.70 = $1.30 per box

Volume in boxes required to earn $1084000 operating profit = (Fixed expense + Target profit) / Contribution per unit

= (440000 +184000) / 1.30

= 480,000 boxes

Volume in boxes required to earn $1084000 operating profit = 480,000 boxes

Answer 4:

Contribution per box of new product = $6.00 - ($3.00 + $1.00) = $2.00

Sales proportion in sales volume in boxes = 90% of couscous + 10% of Taco couscous

Composite contribution of sales mix = 90% * $1.30 + 10% * $2.00 = $1.37

Volume in boxes required to earn $1084000 operating profit

= (440000 +184000) / 1.37

= 455474.45 boxes

Reason for difference in boxes:

We observe, with inclusion of new product, number of boxes required now to earn same operating profit is lesser than number of boxes required (without new product).The reason is contribution per box of new product is higher. Which is resulting in reduction in required number of boxes to same operating profit.